OpenStudy (emac):

Find where it is increasing/decreasing (5x-4)/(x+4)

OpenStudy (emac):

I already found that there were no critical points and I'm not sure how to find where it's increasing or decreasing without any critical points

OpenStudy (mathmale):

If the graph "rises" as x increases, the function is increasing. Does that help? A function that is increasing certainly can increase without critical points as markers. The CP do define where the function begins to increase (or decrease), however.

OpenStudy (emac):

Would that mean a positive function is increasing from (-infinity, infinity)?

OpenStudy (irishboy123):

\(y = \dfrac{5x-4}{x+4} = \dfrac{5(x+4) - 20-4}{x+4} \) \(= 5 - \dfrac{24}{x+4} \) \(Y = - \dfrac{24}{x+4} \) where \(Y = y - 5\) \(Y = - \dfrac{24}{X} \) where \(X = x + 4\) and so on :)

OpenStudy (irishboy123):

in terms of critical or fixed points, we have just made linear transformations, ...and find that \(\dfrac{d Y }{dX} = \dfrac{24}{X^2}\) it follows too that \(\dfrac{d X }{dY} = \dfrac{X^2}{24}\)

OpenStudy (emac):

Where did you get 5(x+4)-20-4 on the top of the fraction?

OpenStudy (irishboy123):

\(y = \dfrac{5x-4}{x+4} = \dfrac{\color{red}{5(x+4) - 20-4}}{x+4}\) \(= \dfrac{\color{red}{5x+20 - 20-4}}{x+4}\) do you disagree?

OpenStudy (emac):

I'm just confused why you changed 5x-4 into 5(x+4)-20-4

OpenStudy (irishboy123):

let me re-think it... if you don't like it

OpenStudy (irishboy123):

yes, right at the start, i remember. it's all about simplifying that expression

OpenStudy (mathmale):

Your function is y=(5x-4)/(x+4), or \[y=\frac{ 5x-4 }{ x+4 }\]

OpenStudy (mathmale):

Jump in and take the derivative immediately:\[\frac{ dy }{ dx }=\frac{ (x+4)(5) - (5x-4)(1) }{ (x+4)^2 }\] Here I used the Quotient Rule. Please be absolutely certain that you understand this; if you don't, ask for clarification. The derivative simpifies to\[\frac{ dy }{ dx }=\frac{ 5x+20-5x+4 }{ (x+2)^2 }=\frac{ 24 }{ (x+2)^2 }\]

OpenStudy (mathmale):

True, you cannot find the critical value(s) by setting this = to 0 and solving for x. HOWEVER, you do have a critical value at x=-2, because the derivative is not defined there. Plot x=-2 on a number line. On the left you'll have (-inf,-2) and on the right you'll have (-2, inf). From each of these 2 intervals, choose a test number. Subst. this test number into the first derivative to determine whether the deriv. is pos. or neg. Know what to do next? If you'd like feedback, please share your work. Thx.